NASA ISS On-Orbit Status 6 August 2013

©NASA

NASA ISS On-Orbit Status 6 August 2013.

New Status: H-II Transfer Vehicle 4 (HTV4) Status: Yesterday HTV4 telemetry indicated that the Reaction Control System (RCS)-A latching valve was closed unexpectedly. Ground teams tested the latching valve today and determined that both the latch and RCS-A are working nominally.

They concluded that the anomaly was a transient in the switch that detects the valve open / close status. Currently RCS-B is prime and RCS-A is in a backup mode. HTV4 is flying nominally and no impacts are expected to the planned capture and berthing operations on Friday. Today HTV executed a planned 13 second burn successfully. Also today Lane 2 of the Service Module (SM) [TVM] navigation computer experienced issues. However with 2 good [TVM] lanes still functioning nominally, there is no impact to HTV rendezvous planned for Friday. Russian specialists are investigating the issue.

Robonaut Operations: Today, Flight Engineer (FE)-3 Cassidy, in cooperation with the Robonaut Principal Investigator (PI), performed Robonaut Tele-operations. Cassidy activated Robonaut and the Robonaut Tele-operation System (RTS) which consists of 3D visor, vest and gloves. These hardware items worn by the operator contain specialized sensors capable of detecting the motion of the operator. The sensors transmit operator movements to a computer that works through the existing Robonaut control software to drive the robot in fashion to mimics the motion of the operator. Feedback from the robot is in the form of left/right video cameras from Robonaut that feed direct video to a video display helmet worn by the operator. The PI guided Cassidy in performing the Audio Terminal Unit (ATU) actuation, having Robonaut mimic that task. Robonaut is a dexterous humanoid robot designed with the versatility and dexterity to manipulate hardware, exhibit greater endurance than humans and react safely when bumped or interacted with in a way that was not expected.

Marangoni Inside (MI) Experiment Operations: FE-6 Nyberg disassembled and stowed the MI Ultrasonic Velocity Profile (UVP)-2R cartridge. The Spatio-temporal Flow Structure in Marangoni Convection (Marangoni-UVP) investigation is one of several investigations carried out by (Japan Aerospace Exploration Agency) JAXA to investigate Marangoni convection, a process driven by the presence of a surface tension gradient as produced by temperature differences occurring at a liquid/gas interface. Oscillatory convection, three-dimensional (3D) flow and particle accumulation structure (PAS) formation are all ideally observed using the Fluid Physics Experiment Facility (FPEF) and the ideal convection characteristics afforded by the microgravity environment. Continued research and understanding of liquid flow characteristics may lead to the development highly efficient industrial process and products involving semiconductors, optical materials, bio materials, thermal transport devices.

Urine Processing Assembly (UPA) Status: Last Thursday, the UPA shut down due to a fault as a result of high pressure in the system. Today FE-5 Parmitano successfully removed and replaced the Pressure Control and Pump Assembly (PCPA) with one of two on orbit spares. Ground Controllers are watching pressure data and the pressure is dropping as expected. Once the pressure is at an acceptable level, a process cycle will be initiated.

Airlock Video Troubleshooting: FE-5 Parmitano continued troubleshooting the Airlock video system and determined that the Audio Visual Interface Unit (AVIU) was failed. He installed a spare AVIU and preliminary analysis indicates the system is functioning nominally. There are 2 AVIUs that are considered suspect and 1 AVIU onboard that is considered good.

Mobile Servicing System (MSS) Operations: Robotics Ground Controllers and Visiting Vehicle Officers (VVO) performed part 1 of the Robotics Workstation (RWS) ground camera calibration in preparation for HTV4 capture. Prior to the start of this procedure, Cassidy set up the LAB Camcorder to provide a view of the LAB RWS Monitor 1. VVO took screenshots of the current pointing of the S1 Lower Outboard (S1LOOB), the P1 Lower Inboard (P1LOIB), the Special Purpose Dexterous Manipulator System (SPDM) Camera Light Pan Tilt Unit (PTU) Assembly (CLPA) 2 and the P1LOOB cameras as displayed on Lab RWS Monitor 1 to assess camera calibration requirements. Upon completion of this procedure, VVO reported that S1LOOB, P1LOIB and P1LOOB did not need to be calibrated. SPDM CLPA2 pointing was slightly off, so VVO will review the data collected to determine if calibration is required. Part 2 of the calibration, if required, will be completed Thursday. MSS performance today was nominal.

Today's Planned Tasks

Reaction Self Test [Complete]
SPRUT-2 Experiment Ops / r/g 3119 [Complete]
Reaction Self Test [Complete]
Laptop RS2(1) Reboot [Complete]
CASKAD. Manual Mixing in Bioreactor / r/g 3070 [Complete]
Set up video equipment to record ARED exercise [Complete]
WRS Water Sample Analysis [Complete]
Pretreat and Water Dispenser (ДкиВ) Replacement [Complete]
ASEPTIC. Samples photo after incubation. r/g 3114 [Complete]
ELT2 Setup [Complete]
CIR - Hardware Setup [Complete]
ROBONAUT - procedure review [Complete]
URAGAN. Photo Spectrum System Battery Charge (start) / r/g 2857 [Complete]
Replace the kit content of charger set, МВП-2К electronics box, and Scope Meter Fluke 105B [Complete]
ASEPTIC. Photo Edit and Downlink r/g 3114 [Complete]
ROBONAUT- Camcorder Setup [Complete]
R2-Hardware Setup and Video [Complete]
Audio Dosimeter Preparation [Complete]
ROBONAUT- Camcorder Setup [Complete]
ASEPTIC. Cryogem-03 Removal / r/g 3115 [Complete]
ROBONAUT Tele-ops Hardware Setup [Complete]
Photo/TV Camcorder Setup Verification [Complete]
Video T/S [Complete]
ROBO - Assistance with the Experiment [Complete]
Robonaut Tele-Ops [In work]
ТКГ 420 Transfers and IMS Updates (r/g 3035)+ r/g 3043 , 3032, 3047
MI-MSPR. Installation of Work Bench [Complete]
ISS - HAM Radio Session [Complete]
Gather EVA Equipment and Tools. Photography and Photo Downlink [Complete]
G1 - Camcorder Setup [Complete]
MI - Hardware Setup [Complete]
Photo/TV Camcorder Setup Verification [Complete]
TOCA 2 Data Recording [Complete]
Health Maintenance System (HMS) Tonometry Test Setup [Complete]
TOCA Waste Water Bag (WWB) Changeout [Complete]
HMS - Tonometry Test - CMO [Complete]
HMS - Tonometry Test - CMO [Complete]
ELT2 Reboot [Complete]
Node2 Camcorder Reconfiguration [Complete]
MI. Core Removal [Complete]
ROBONAUT - Reconfigure Camcorder for Nominal Ops [Complete]
Stow after Tonometry Test [Complete]
ROBONAUT - Hardware Disassembly [Complete]
Gather EVA Equipment and Tools. Photography and Photo Downlink [Complete]
URAGAN. Hardware Stowage [Complete]
Pressure Control and Pump Assembly [PCPA] R&R in WRS2 [Complete]
Marangoni Inside (MI) Interface Stow [Complete]
Photo/TV Camcorder Setup Verification [Complete]
Hand Ergometry Assesment.
G1 - Closeout Ops [Complete]
MI-MSPR. Close-out ops with work bench [Complete]
Gather EVA Equipment and Tools [Complete]
PAO Hardware Setup [Complete]
Crew Prep for PAO [Complete]
Pressure Control and Pump Assembly Restow [Complete]
PAO Event [Complete]
EVA Mag Test [Complete]
Hand Ergometry. Tagup with specialists (VHF) [Complete]
ТКГ 420 Transfers and IMS Updates + r/g 3043 , 3032, 3047
PMM Clean up Followup [Complete]
IMS Update
CASKAD. Manual Mixing in Bioreactor / r/g 3070
RWS - Video Downlink
Reaction Self Test
Reaction Self Test
Task List

ARED Cyl Evac [Complete]
WHC KTO Replace
WHC EDV R&R [Complete]
Roscosmos Report
52P USOS Unpack [In Work]
Bk Curtain Retrieve
ARED Detent Grease [Complete]
Data Prep for Return FE-3
TVIS Cuecard Discard
Regen WSTA Fill
Ground

Nominal Commanding
Three-Day Look Ahead:

Wednesday, 8/07: Spheres, ACE, Ultrasound, BCAT

Thursday, 8/09: CFE, HTV Robot Training, BCAT, Fundoscope

Friday, 8/09: HTV Capture & Berthing, INSPACE3, HTV Vestibule Press

QUICK ISS Status - Environmental Control Group:

Please follow SpaceRef on Twitter and Like us on Facebook.